# Scientific Machine Learning of Chaotic Systems Discovers Governing Equations for Neural Populations

### v2

@@ II.1.4 Combining PEM-UDE into single method
$\frac{d\tilde{\mathbf{x}}}{dt}$ $=f_{p}(\tilde{\mathbf{x}},\mathbf{u}(t))+K\epsilon_{\theta}(t)+NN_{\theta}(\tilde{\mathbf{x}})$ (6a)
@@
In this work, we consider only systems where we have both an initial functional form and observational data, or at least an initial functional form, as the estimation of unobserved latent states for which we do not have an initial idea of the functional form is a separate problem that has been discussed elsewhere [43, 34]. As we are focused on the training of UDE systems with PEM – an inherently data-driven approach – it is beyond the scope of the current work to focus on these approaches, which could be combined in future work on more complicated systems that include both explicit and latent dynamics. The results are organized to progress from high information systems where we can both form a strong initial hypothesis of the dynamics and observe all the states to progressively less information both of the functional form and the observability of the states.
@@ II.1.5 Symbolic recovery of terms from PEM-trained UDE
After fitting the dynamics of the system with this combined PEM-UDE approach, we then extract a symbolic form of the dynamics from the universal approximator $NN_{\theta}$. This typically provides a system that generalizes to future contexts more robustly than a “black-box” neural network [16, 3], and allows for direct analysis of the learned functional form. We demonstrate here that the terms learned by the UDEs can be recovered by either sequential thresholding least squares (STLSQ), which learns features from a library via ridge regression [10], or via a genetic algorithm approach to symbolic regression [14]. Each approach has its particular strengths and weaknesses, and selecting the correct approach will depend on the structure of the data and known dynamics. Genetic algorithm symbolic regression offers a better approach where there are complex constraints on the functional form of the system [14], while STLSQ works better on systems that require very many data points to achieve an accurate fit. We offer both approaches in the code accompanying this method, and discuss our selection of symbolic recovery method for each the sections individually below. As these approaches are not novel to this work, we save further discussion of them for the Methods.
@@ II.2 PEM-UDE outperforms UDE in fitting chaotic systems
We begin by demonstrating the utility of the PEM-UDE approach in fitting the Rössler system (Fig. 2). This system was chosen as it is a simple set of ODEs with only a single nonlinear term and yet has a rich set of chaotic dynamics for the right parameter set [22]. In this demonstration, we assume we know the functional form of $x$ and $z$ but have no information about the functional form of $y$, and so must learn the dynamics from the time series observations. With only this challenge, a traditional UDE fails to fit well (Fig. 2a), with the UDE-learned dynamics typically converging to the mean of the time series (typically a local minimum in the objective function). By contrast, the PEM-UDE approach is able to converge successfully after training, with a clear difference in performance even after the first 10% of the total training time (at which point the standard UDE has already neared its local minimum).
@@
Figure 2: The combined PEM-UDE approach is able to learn the dynamics of chaotic systems that traditional UDEs are unable to capture. a. Comparison of training error for the UDE system (fails to accurately capture dynamics) to the PEM-trained UDE (succeeds). The difference in accuracy is already stark 10% into the full training session (inset), where unassisted UDEs are . b. The application of PEM smooths out an otherwise intractable parameter loss landscape, with the hyperparameter $K$ tuning the steepness of the landscape. c. Comparison of the error of the fit dynamics of the UDE to the PEM-trained UDE. d. Using STLSQ to replace the PEM-trained UDE results in an expression that not only captures the original dynamics (first 200 time points) but continues to accurately describe them well into the future (last 100 time points). e. The surface of the Rössler attractor demonstrates chaotic filling of phase space, with a basin near the x-y plane and occasional bursts along the z-axis. f. Relative error of the PEM-trained UDE at each point of the attractor surface.
@@
The reason for this performance increase lies in the way that the PEM approach alters the landscape over which the UDE is being trained (Fig. 2b). We note that the unmodified landscape is extremely rough and contains several local minima, with only a very sharp decrease in loss at the true global minimum ($a=0.2$ in this case). By adding the PEM to the dynamics of the system during training of the UDE, the landscape becomes much more approachable via standard optimization techniques. Further, the choice of $K$ can be made to directly tune the steepness of the landscape. This should be done in conjunction with considering how the PEM will smooth out the landscapes of other parameters in states where the PEM is not applied. For reasonable choices of $K$, the addition of PEM will not substantially deviate the fit dynamics from the true minima along other parameters (explored more fully in Fig. 6). Importantly, as higher levels of filtering are introduced with larger $K$, finding the true functional form of the entire system will rely more upon the accuracy of the estimates of the un-fit parameters (Fig. 7). When using the PEM-UDE combination approach, it is therefore best to select a hyperparameter $K$ that provides a smooth enough landscape for optimization to occur but is steep enough that the global minimum is still clearly distinguishable from the surrounding area to allow for noisier estimates of unmeasured parameters.
@@
At its core, the PEM-UDE approach works because the prediction error correction removes the sensitive dependence on parameters from the system being fit, effectively negating the hallmark trait of chaotic systems. As shown in Fig. 9a, when $K$ is small there is an $\mathcal{O}(1)$ divergence with respect to any deviation in the initial parameter choice (illustrated as a deviation of 1e-9 in Fig. 9a), but as $K$ is sufficiently large ($K\approx 0.18$ in this example) this sensitive dependence is eliminated and the error between the two trajectories shrinks. For sufficiently large $K$, this error will depend on the size of the deviation in the parameter, and $K$ can be tuned based on the estimate of this deviation (Fig. 9b). This property, along with the fact that when the error term in Eq. 6a shrinks to zero the dynamics of the original system are fit, means that the PEM perturbed system is well-behaved (i.e. non-chaotic) and has a loss minimum at the same point as the original system. As such, the L2 norm trajectory fitting approach which is unstable on the original chaotic system is thus stabilized on the PEM-UDE, allowing the fitting process to convergence while leading to an effectively unbiased final fit.
@@
To generalize these dynamics beyond the original observations, we extracted a symbolic form of the equations using STLSQ fit to the dynamics of the PEM-UDE system and the outputs of the UDE at each timestep. The functional form of the missing state is directly recovered by STLSQ ($y=p_{1}x+p_{2}y$), with the initial fits of parameters close to the true values ($p_{1}=1.01$, $p_{2}=0.17$, while the true values are $p_{1}=1$, $p_{2}=0.2$). By adding an additional optimization step to tune the parameters of this learned symbolic equation after STLSQ, we are able to converge to the true values (accurate to three decimal places), which then generalize the dynamics forward in time beyond the original simulations (Fig. 2d). Importantly, this pipeline provides an analytical form of the dynamics that not only generalizes forwards in time but can also be interrogated for insights into the physical system.
@@ II.3 PEM-UDE can learn chaotic dynamics even if certain states have limited observational data
Figure 3: The PEM-UDE approach is capable of recovering the correct form of a chaotic circuit even in the presence of significant observational noise. a. The analog Petrzela-Polak circuit that produces the dynamics examined in this section. The observed states are the dimensionless forms of $V_{1}$, $V_{2}$, and $I$, where $V_{2}$ and $I$ (green boxes) are observed with only a small amount of noise, and $V_{1}$ (blue box) is observed with large observational noise. b. In this example system, we assume a faulty observation of one states ($V_{1}/z$), where the true dynamics (top) are masked by observational noise of 5x the true signal’s magnitude (bottom). c. Original attractor surface, with true form of the observed dimension $x$ in color with magnitude of error from observational noise shown. d. Attractor surface that is fit by the PEM-UDE method with noise in the third dimension (notice the scattered colors and the noise has significantly corrupted the dynamics of the state $x$). e. The true form of the state to be recovered is shown at the top, with the form recovered by STLSQ applied to the PEM-UDE dynamics is shown at the bottom. The functional form - although not the exact parameters - is recoverd from the PEM-UDE fit (table compares parameters fit by STLSQ to original parameters).
We next present a second example of the PEM-UDE method being fit to a chaotic system, with two important differences compared to the Rössler system. First, this example is one of a physically realized electrical circuit, with the circuit diagram shown in Fig. 3a, and the parameters for the physical circuit (as described in [53]) given in 2. Conceptually, this offers a transition from the idealized chaotic dynamics of the previous system to more real-world applications, as observing chaos in electrical circuits has practical implications [54]. To extend our method further, we create a second challenge – adding a noisy observation to one of the circuit components ($V_{1}$, denoted by the red box shown in Fig. 3a). The true equations for this system are given in 14a, with the dimensionless form used in this example given by 15a. In this section, we assume there is an unknown faulty observation of the first state $x$, such that there is noise around the observation at 5x the magnitude of the true signal (shown in Fig. 3b). This creates a difficulty for the PEM-UDE fit: the true dynamics (Fig. 3c) are now masked by a significant amount of noise that grossly distorts the shape of the attractor being fit (Fig. 3d).
@@ 0.5 $\phi_{1}+\phi_{2}y+\phi_{3}z$
We assume that this noise is unknown at the time of fitting and proceed with fitting the PEM-UDE to the corrupted data using the same approach as above (i.e., fitting the functional form shown in 16a). Despite the significant observational noise present in one of the states, STLSQ applied to the PEM-UDE fit dynamics recovers the exact functional form of the dynamics (Fig. 3e). However, we note that this approach does not provide the original parameters, although the form is correct. In this approach, however, we note that we have shifted the problem from one of unknown dynamics to one of simple parameter optimization, to which other approaches (e.g., simulation-based inference [55]) can be fruitfully applied. This also gives us confidence that even in cases of extremely noisy observational data, the PEM-UDE approach can recover valuable information about the true dynamics of the system.
This example also provides a useful demonstration of why fitting the symbolic form of a trained UDE outperforms a STLSQ fit of the original dataset. As a comparison to the dynamics shown in Fig. 3e, we fit the same noisy dataset directly using the SINDy approach [10, 56], i.e. calculating a dataset of smoothed derivative estimates and performing the STLSQ symbolic regression method directly on the derivative estimates. As shown in Table 1, no recovery of the true symbolic form is possible with this method across a range of the ADMM hyperparameter $\lambda$, with more sparse representations incorrectly dropping the noisy state $x$ from the expression entirely. Recovering the symbolic form of the trained UDE is more accurate as the UDE provides a noise-free (or greatly noise-reduced) version of the dynamics that can be sampled with arbitrary density, thereby mitigating the sensitivity that symbolic regression methods show to noisy data.
@@ II.4 Learning novel expressions for sparse networks that violate original derivation assumptions
Having demonstrated the utility of this approach in learning forms of chaotic physical systems, we next turn to deriving novel expressions for the mean-field activity of neuron populations. The data needed to study the human brain inherently bridges multiple scales, as individual neuronal activity gives rise to synchronized actions that define interacting functional circuits (Fig. 4a-c). Traditional neural mass models represent the average activity of neuron populations by modeling specific phenomena (e.g., excitatory/inhibitory balance in the Wilson-Cowan model [39], ion gradient dynamics in excitatory populations in the Larter-Breakspear model [41, 57]), but lack a direct tie to many biophysical parameters (e.g., neurotransmitters) that are essential in understanding inter-neuron dynamics. To overcome this, next-generation neural mass models (NGNMMs) take a different approach, providing an analytical derivation of the mean-field dynamics that retains the biophysical details of individual neurons while also capturing the emergent properties at the population level [42, 43]. The derivation and its key assumptions are discussed further in the Methods.
@@
This analytical form of mean-field dynamics offered by NGNMMs provides many advantages over traditional neural mass models. First, there is a direct connection between the single neuron parameters and the dynamics of the neural population, a feature that is only available in certain populations of traditional neural masses tuned to be sensitive to these parameters [41, 57]. Second, NGNMMs capture emergent properties of neural populations that traditional neural masses do not. For example, $\beta$-rebound – where the suppressed activity in the $\beta$ band increases in power after being suppressed (a commonly observed phenomenon in motor cortex studies) – is observed as a feature of coupled NGNMMs that are built only to oscillate in the $\beta$ band, with no functional constraints that force this activity to occur [44].
@@
Despite these advantages, however, there are some notable limitations to the NGNMM approach, mostly involving the assumptions necessary for the Ott-Antonsen ansatz [45], a key component of the NGNMM derivation. Most important from a biophysical standpoint is the assumption of all-to-all connectivity of the neuron population. This is approximately correct in deep regions (e.g., hippocampus, entorhinal cortex) that have significant within-region collaterals [43]. Immediate neighbors of these regions tend to have fewer collaterals, however, like the thalamostriatal connection patterns [58]. The assumption of full connectivity becomes obviously false in the cortex [48], which typically has very sparse connectivity (on the order of 1-5% of all possible connections; Fig. 4d). Accounting for sparsity in NGNMMs is, therefore, a key problem that must be solved for them to be generalizable in new contexts.
@@
As there is disagreement on the correct analytical approach to adjust the derivation of NGNMMs for sparsity [47, 59, 60], we propose to sidestep the issue of an analytical derivation entirely and simply learn the form of the sparse networks directly from the data with the PEM-UDE method. Following the analytical approach for Izhikevich neurons [43], we begin with a large population of neurons described by Eq. 17 with synaptic dynamics in Eq. 19. In the all-to-all case, the analytical NGNMM describing the activity of this population is given by Eq. 21a. As we begin to prune connections in the graph to introduce sparsity, however, we see that this form no longer accurately represents the population activity. We therefore learn correction terms for the mean firing rate $r$ and voltage $v$ given a reduced probability of connectivity $p(c)$, where $p(c)\in[0.05,1]$. Although there are additional states for the recovery current $w$ and the synaptic dynamics $s$, we only learn from the firing rate and voltage states and train only on their data, as these are the data typically available from experiments. Learning directly from these data sources thus allows us to demonstrate the applicability of this approach in learning directly from experimental data.
@@
The PEM-UDE approach is able to capture the mean-field dynamics of the network under different degrees of sparsity to a high degree of accuracy (Fig. 4e). This accuracy does not include the initial conditions; this is discussed further in the limitations and future directions portion of the discussion. It does, however, capture a noticeable shift in the dominant frequency of oscillations present in the network as a function of sparsity (Fig. 4f). Densely connected regions typically have a primary frequency peak in the $\theta$-band, while the most sparse networks exhibit a primary peak in the $\alpha$-band. This emerges as a natural feature of the learned dynamics and is also observed in experimental data. Hippocampal rhythms are typically in $\theta$ frequencies and emerge from the densely connected areas of the brain [61], while $\alpha$ frequencies typically emerge from sparsely connected cortical regions before propagating through the brain [62]. The PEM-UDE learned dynamics therefore preserve and extend the strength of NGNMMs in capturing the emergent properties (e.g., synchrony) at the mean-field scale by accurately captuging the ergodic properties of the system.
@@
As in the previous two systems, we extend the PEM-UDE learned dynamics described in Eq. 23a by fitting a symbolic form of the dynamics to the UDE network with STLSQ, providing an analytical form of the solution (shown in Eq. 24a, with fit parameters in Table 4). As NGNMMs are derived based on the assumption of a population of coupled oscillators, there is an exact expression for the degree of synchrony within the population given by the expression for the Kuramoto order parameter $Z$
@@
In the case of the symbolic form of the PEM-UDE learned network dynamics, we observe an emergent prediction that more densely connected regions will tend to have higher degrees of synchrony as a baseline state. In comparison, more sparsely connected regions will tend to exhibit lower degrees of synchrony when not acted upon by an external drive (Fig. 4g). This forms an interesting and unexpected prediction from the discovered equations for networks of different sparsity. To test the validity of this prediction, we examined three different datasets of intracranial electrode recordings from regions of different degrees of sparsity (Fig. 5). These datasets included recordings from the rat medial entorhinal cortex (MEC; N=6), a deep region near the hippocampus with high intraregional connectivity [63, 64]; the rat prefrontal cortex (N=10), which has significantly sparser connectivity [65, 66]; and human orbitofrontal cortex (N=10), which exhibits a similarly high degree of sparsity [67, 68]. As predicted by the discovered equations, we note that the deeper, more densely connected structures exhibit a higher degree of synchrony than either of the cortical datasets, and this is a broad trend across individuals, although there is significant variability between subjects. These results suggest that the learned equations are not merely capturing the dynamics on which they are trained but have been able to predict features beyond their initial training data, indicating a more general applicability than the original form allowed.

### v3

@@ II.1.4 Combining PEM-UDE into single method
[…]
@@
In this work, we consider only systems where we have both an initial functional form and observational data, or at least an initial functional form. Problems of estimating unobserved latent states without an initial functional form have been discussed elsewhere [chen2022, antal2024]. As we focus on training UDE systems with PEM – an inherently data-driven approach – it is beyond the scope of the current work to address more complex systems that include both explicit and latent dynamics. Our results are organized to progress from high-information systems, where we can both form a strong initial hypothesis about the dynamics and observe all states, to progressively lower information, both about the functional form and the observability of the states.
@@ II.1.5 Symbolic recovery of terms from PEM-trained UDE
After fitting the dynamics of the system with this combined PEM-UDE approach, we then extract a symbolic form of the dynamics from the universal approximator $NN_{\theta}$. This typically provides a system that generalizes to future contexts more robustly than a “black-box” neural network [rackauckas2021, cranmer2020], and allows for direct analysis of the learned functional form. We demonstrate here that the terms learned by the UDEs can be recovered by either sequential thresholding least squares (STLSQ), which learns features from a library via ridge regression [brunton2016], or via a genetic algorithm approach to symbolic regression [cranmer2023]. Each approach has its particular strengths and weaknesses, and selecting the correct approach will depend on the structure of the data and known dynamics. The symbolic regression using a genetic algorithm offers a better approach when there are complex constraints on the functional form of the system [cranmer2023], while STLSQ works better on systems that require large numbers of data points to achieve an accurate fit. We offer both approaches in the code accompanying this method, and discuss our selection of symbolic recovery method for each of the sections individually below. As these approaches are not novel to this work, we save further discussion of them for the Methods.
@@ II.2 PEM-UDE outperforms UDE in fitting chaotic systems
We begin by demonstrating the utility of the PEM-UDE approach in fitting the Rössler system (Fig. 2). This system was chosen as it is a simple set of ODEs with only a single nonlinear term and yet has a rich set of chaotic dynamics for the right parameter set [rossler1976]. In this demonstration, we assume we know the functional form of $x$ and $z$ but have no information about the functional form of $y$, and so must learn the dynamics from the time series observations. With only this challenge, a traditional UDE fails to fit well (Fig. 2a), with the UDE-learned dynamics typically converging to the mean of the time series (typically a local minimum in the objective function). By contrast, the PEM-UDE approach is able to converge successfully after training, with a clear difference in performance even after the first 10% of the total training time (at which point the standard UDE has already neared its local minimum).
@@
Figure 2: The combined PEM-UDE approach is able to learn the dynamics of chaotic systems that traditional UDEs are unable to capture. a. Comparison of training error for the UDE system (fails to accurately capture dynamics) to the PEM-trained UDE (succeeds). The difference in accuracy is already stark 10% into the full training session (inset), where unassisted UDEs are not progressing. b. The application of PEM smooths out an otherwise intractable parameter loss landscape, with the hyperparameter $K$ tuning the steepness of the landscape. c. Comparison of the mean-square-error of the fit dynamics of the UDE to the PEM-trained UDE. d. Using STLSQ to replace the PEM-trained UDE results in an expression that not only captures the original dynamics (first 200 time points) but continues to accurately describe them well into the future (last 100 time points). e. The surface of the Rössler attractor demonstrates chaotic filling of phase space, with a basin near the x-y plane and occasional bursts along the z-axis. f. Relative error of the PEM-trained UDE at each point of the attractor surface.
@@
The reason for this performance increase lies in the way that the PEM approach alters the landscape over which the UDE is trained (Fig. 2b). We note that the unmodified landscape is extremely rough and contains several local minima, with only a very sharp decrease in loss at the true global minimum ($a=0.2$ in this case). By adding the PEM to the dynamics of the system during training of the UDE, the landscape becomes much more approachable via standard optimization techniques. Further, the choice of $K$ can be used to tune the steepness of the landscape directly. This should be done in conjunction with considering how the PEM will smooth out the landscapes of other parameters in states where the it is not applied. For reasonable choices of $K$, the addition of PEM will not substantially deviate the fit dynamics from the true minima along other parameters (explored more fully in Fig. 6). Importantly, as higher levels of filtering are introduced with larger $K$, finding the true functional form of the entire system will rely more upon the accuracy of the estimates of the un-fit parameters (Fig. 7). When using the PEM-UDE combination approach, it is therefore best to select a hyperparameter $K$ that provides a smooth enough landscape for optimization to occur but is steep enough that the global minimum is still clearly distinguishable from the surrounding area to allow for noisier estimates of unmeasured parameters.
@@
At its core, the PEM-UDE approach works because the prediction error correction removes the sensitive dependence on parameters from the system being fit, effectively negating the hallmark trait of chaotic systems. As shown in Fig. 9a, when $K$ is small there is an $\mathcal{O}(1)$ divergence with respect to any deviation in the initial parameter choice (illustrated as a deviation of $10^{-9}$ in Fig. 9a), but as $K$ is sufficiently large ($K\approx 0.18$ in this example) this sensitive dependence is eliminated and the error between the two trajectories shrinks. For sufficiently large $K$, this error will depend on the size of the deviation in the parameter, and $K$ can be tuned based on the estimate of this deviation (Fig. 9b). This property, along with the fact that when the error term in Eq. 6a shrinks to zero as the dynamics of the original system are fit, means that the PEM perturbed system is well-behaved (i.e. non-chaotic) and has a loss minimum at the same point as the original systeAs such, the L2 norm trajectory fitting approach, which is unstable on the original chaotic system, is thus stabilized on the PEM-UDE, allowing the fitting process to converge while leading to an effectively unbiased final fit.it.
@@
To generalize these dynamics beyond the original observations, we extracted a symbolic form of the equations using an STLSQ fit to the outputs of the UDE taken at each time step of the dynamics of the PEM-UDE system. The functional form of the missing state is directly recovered by STLSQ ($y=p_{1}x+p_{2}y$), with the initial fits of parameters close to the true values ($p_{1}=1.01$, $p_{2}=0.17$, while the true values are $p_{1}=1$, $p_{2}=0.2$). By adding an additional optimization step to tune the parameters of this learned symbolic equation after STLSQ, we are able to converge to the true values (accurate to three decimal places), which then generalize the dynamics forward in time beyond the original simulations (Fig. 2d). Importantly, this pipeline provides an analytical form of the dynamics that not only generalizes forwards in time but can also be interrogated for insights into the physical system.
@@ II.3 PEM-UDE can learn chaotic dynamics even if certain states have limited observational data
Figure 3: The PEM-UDE approach is capable of recovering the correct form of a chaotic circuit even in the presence of significant observational noise. a. The analog Petrzela-Polak circuit that produces the dynamics examined in this section. The observed states are the dimensionless forms of $V_{1}$, $V_{2}$, and $I$, where $V_{2}$ and II (green boxes) are observed with only a small amount of noise, and $V_{1}$ (blue box) is observed with large observational noise. b. In this example system, we assume a faulty observation of one state ($V_{1}$), where the true dynamics (top) are masked by observational noise of 5x the true signal’s magnitude (bottom). c. Original attractor surface, with true form of the observed dimension $x$ in color with magnitude of error from observational noise shown. d. Attractor surface that is fit by the PEM-UDE method with noise in the third dimension (notice the scattered colors and the noise has significantly corrupted the dynamics of the state $x$). e. The true form of the state to be recovered is shown at the top, with the form recovered by STLSQ applied to the PEM-UDE dynamics is shown at the bottom. The functional form - although not the exact parameters - is recoverd from the PEM-UDE fit (table compares parameters fit by STLSQ to original parameters).
We next present a second example of the PEM-UDE method being fit to a chaotic system, with two important differences compared to the Rössler system. First, this example is one of a physically realized electrical circuit, with the circuit diagram shown in Fig. 3a, and the parameters for the physical circuit (as described in [sprott2022]) given in Table 2. Conceptually, this offers a transition from the idealized chaotic dynamics of the previous system to more real-world applications, as observing chaos in electrical circuits has practical implications [kuznetsov2023]. To extend our method further, we create a second challenge – adding a noisy observation to one of the circuit components ($V_{1}$, denoted by the red box shown in Fig. 3a). The true equations for this system are given in Eqs. 14, with the dimensionless form used in this example given by Eqs. 15. In this section, we assume there is an unknown faulty observation of the first state $x$, such that there is noise around the observation at 5x the magnitude of the true signal (shown in Fig. 3b). This creates a difficulty for the PEM-UDE fit: the true dynamics (Fig. 3c) are now masked by a significant amount of noise that grossly distorts the shape of the attractor being fit (Fig. 3d).
@@ 0.5 $\phi_{1}+\phi_{2}y+\phi_{3}z$
We assume that this noise is unknown at the time of fitting and proceed with fitting the PEM-UDE to the corrupted data using the same approach as above (i.e., fitting the functional form shown in Eqs. 16). Despite the significant observational noise present in one of the states, STLSQ applied to the PEM-UDE fit dynamics recovers the exact functional form of the dynamics (Fig. 3e). However, we note that this approach does not provide the original parameters, although the form is correct. In this approach, however, we note that we have shifted the problem from one of unknown dynamics to one of simple parameter optimization, to which other approaches (e.g., simulation-based inference [goncalves2020]) can be fruitfully applied. This also gives us confidence that even in cases of extremely noisy observational data, the PEM-UDE approach can recover valuable information about the true dynamics of the system.
This example also provides a useful demonstration of why fitting the symbolic form of a trained UDE outperforms a STLSQ fit of the original dataset. As a comparison to the dynamics shown in Fig. 3e, we fit the same noisy dataset directly using the SINDy approach [brunton2016, juliusmartensen2021], i.e. calculating a dataset of smoothed derivative estimates and performing the STLSQ symbolic regression method directly on the derivative estimates. As shown in Table 1, no recovery of the true symbolic form is possible with this method across a range of the ADMM hyperparameter $\lambda$, with more sparse representations incorrectly dropping the noisy state $x$ from the expression entirely. Recovering the symbolic form of the trained UDE is more accurate as the UDE provides a noise-free (or greatly noise-reduced) version of the dynamics that can be sampled with arbitrary density, thereby mitigating the sensitivity that symbolic regression methods show to noisy data.
@@ II.4 Learning novel expressions for sparse networks that violate original derivation assumptions
Having demonstrated the utility of this approach in learning forms of chaotic physical systems, we next turn to deriving novel expressions for the mean-field activity of neuron populations. The data needed to study the human brain inherently bridges multiple scales, as individual neuronal activity gives rise to synchronized actions that define interacting functional circuits (Fig. 4a-c). Traditional neural mass models represent the average activity of neuron populations by modeling specific phenomena (e.g., excitatory/inhibitory balance in the Wilson-Cowan model [wilson1972], ion gradient dynamics in excitatory populations in the Larter-Breakspear model [larter1999, breakspear2003]), but lack a direct tie to many biophysical parameters (e.g., neurotransmitters) that are essential in understanding inter-neuron dynamics. To overcome this, next-generation neural mass models (NGNMMs) take a different approach, providing an analytical derivation of the mean-field dynamics that retains the biophysical details of individual neurons while also capturing the emergent properties at the population level [montbrio2015, chen2022]. The derivation and its key assumptions are discussed further in the Methods.
@@
This analytical form of mean-field dynamics offered by NGNMMs provides many advantages over traditional neural mass models. First, there is a direct connection between the single neuron parameters and the dynamics of the neural population, a feature that is only available in certain populations of traditional neural masses tuned to be sensitive to these parameters [larter1999, breakspear2003]. Second, NGNMMs capture emergent properties of neural populations that traditional neural masses do not. For example, $\beta$-rebound – where the suppressed activity in the $\beta$ band increases in power after being suppressed (a commonly observed phenomenon in motor cortex studies) – is observed as a feature of coupled NGNMMs that are built only to oscillate in the $\beta$ band, with no functional constraints that force this activity to occur [coombes2019].
@@
Despite these advantages, however, there are some notable limitations to the NGNMM approach, mostly involving the assumptions necessary for the Ott-Antonsen ansatz [ott2008a], a key component of the NGNMM derivation. Most important from a biophysical standpoint is the assumption of all-to-all connectivity of the neuron population. This is approximately correct in deep regions (e.g., hippocampus, entorhinal cortex) that have significant within-region collaterals [chen2022]. Immediate neighbors of these regions tend to have fewer collaterals, however, like the thalamostriatal connection patterns [ding2010]. The assumption of full connectivity becomes obviously false in the cortex [pathak2024], which typically has very sparse connectivity (on the order of 1-5% of all possible connections; Fig. 4d). Accounting for sparsity in NGNMMs is, therefore, a key problem that must be solved for them to be generalizable in new contexts.
@@
As there is disagreement on the correct analytical approach to adjust the derivation of NGNMMs for sparsity [divolo2018, goldobin2021, divolo2022, clusella2024], we propose to sidestep the issue of an analytical derivation entirely and simply learn the form of the sparse networks directly from the data with the PEM-UDE method. Following the analytical approach for Izhikevich neurons [chen2022], we begin with a large population of neurons described by Eqs. 17 with synaptic dynamics in Eqs. 19. In the all-to-all case, the analytical NGNMM describing the activity of this population is given by Eqs. 20. As we begin to prune connections in the graph to introduce sparsity, however, we see that this form no longer accurately represents the population activity. We therefore learn correction terms for the mean firing rate $r$ and voltage $v$ given a reduced probability of connectivity $p(c)$, where $p(c)\in[0.05,1]$. Although there are additional states for the recovery current $w$ and the synaptic dynamics $s$, we only learn from the firing rate and voltage states and train only on their data, as these are the data typically available from experiments. Learning directly from these data sources thus allows us to demonstrate the applicability of this approach in learning directly from experimental data.
@@
The PEM-UDE approach is able to capture the mean-field dynamics of the network under different degrees of sparsity to a high degree of accuracy (Fig. 4e). This accuracy does not include the initial conditions; this is discussed further in the limitations and future directions portion of the discussion. It does, however, capture a noticeable shift in the dominant frequency of oscillations present in the network as a function of sparsity (Fig. 4f). Densely connected regions typically have a primary frequency peak in the $\theta$-band, while the most sparse networks exhibit a primary peak in the $\alpha$-band. This emerges as a natural feature of the learned dynamics and is also observed in experimental data. Hippocampal rhythms are typically in $\theta$ frequencies and emerge from the densely connected areas of the brain [lubenov2009], while $\alpha$ frequencies typically emerge from sparsely connected cortical regions before propagating through the brain [halgren2019]. The PEM-UDE learned dynamics therefore preserve and extend the strength of NGNMMs in capturing the emergent properties (e.g., synchrony) at the mean-field scale by accurately captuging the ergodic properties of the system.
@@
As in the previous two systems, we extend the PEM-UDE learned dynamics described in Eqs. 22 by fitting a symbolic form of the dynamics to the UDE network with STLSQ, providing an analytical form of the solution (shown in Eqs. 23, with fit parameters in Table 4). As NGNMMs are derived based on the assumption of a population of coupled oscillators, there is an exact expression for the degree of synchrony within the population given by the expression for the Kuramoto order parameter $Z$
@@
In the case of the form of the PEM-UDE learned network dynamics, we observe an emergent prediction that more densely connected regions will tend to have higher degrees of synchrony as a baseline state. In comparison, more sparsely connected regions will tend to exhibit lower degrees of synchrony when not acted upon by an external drive (Fig. 4g). This forms an interesting and unexpected prediction from the discovered equations for networks of different sparsity. To test the validity of this prediction, we examined three different datasets of intracranial electrode recordings from regions of different degrees of sparsity (Fig. 5). These datasets included recordings from the rat medial entorhinal cortex (MEC; N=6), a deep region near the hippocampus with high intraregional connectivity [hernandez-perez2020, hernandez-perez2020a]; the rat prefrontal cortex (N=10), which has significantly sparser connectivity [peyrache2009, peyrache2018]; and human orbitofrontal cortex (N=10), which exhibits a similarly high degree of sparsity [saez2018, saez2018a]. As predicted by the discovered equations, we note that the deeper, more densely connected structures exhibit a higher degree of synchrony than either of the cortical datasets, and this is a broad trend across individuals, although there is significant variability between subjects. These results suggest that the learned equations are not merely capturing the dynamics on which they are trained but have been able to predict features beyond their initial training data, indicating a more general applicability than the original form allowed.